A spherical shell a diameter of 20 cm it is given charge 1 c. What is the potential at the surface of the shell

  • A spherical shell a diameter of 20 cm it is given charge 1 c.                     What is the potential at the surface of the shell

Grade:12

1 Answers

Khimraj
5 years ago
At surface
V = kq/R = 9*109*1/0.1 = 9*1010 V
Hope it clears.............................................................................